# Instruction-Tuning Pretrained Causal Language Models for Text Restoration, Chronological Attribution, and Geographic Attribution of Ancient Greek Papyri and Inscriptions

### v2

@@ Abstract
This article presents an experiment in fine-tuning a pretrained causal language model (Meta’s Llama 3.1 8B Instruct) to assist with three key tasks in philological research: dating, localizing, and restoring missing or illegible characters in ancient Greek inscriptions and documentary papyri.
Utilizing a straightforward instruction-based approach and a 95%/5% train/test split, the papyri models achieved a character error rate (CER) of 14.9%, a top-1 accuracy of 73.5%, and a top-20 accuracy of 86.0% in text reconstruction. For geographic attribution, they reached a top-1 accuracy of 66.4% and a top-3 accuracy of 79.9%. In chronological attribution, the models demonstrated an average deviation of 21.7 years from the actual terminus post/ante quem, with a median deviation of 0 years.
For inscriptions, the models achieved a CER of 20.5%, a top-1 accuracy of 63.7%, and a top-20 accuracy of 83.0% for sequences up to 10 characters. In geographic attribution, they attained a top-1 accuracy of 75.0% and a top-3 accuracy of 83.7%, while in dating, they had an average deviation of 37.1 years and a median deviation of 3 years from the actual date range.
Benchmarked against the state-of-the-art model (Ithaca) on a shared test set and on recently edited inscriptions, the instruction-tuned models excelled in text restoration, while also offering the practical advantage of ignoring spaces during reconstruction, which aligns with the scriptio continua format of ancient texts. However, their performance in geographic and chronological attribution was lower than Ithaca’s.
The results are preliminary but suggest that fine-tuning larger pretrained causal language models using instruction templates for philological research holds significant promise, particularly in the area of textual criticism.
@@ 1 Introduction
A significant number of documentary texts in ancient Greek – such as announcements, laws, receipts, contracts, and letters – have survived from the Archaic period through to the early Byzantine era. These documents were inscribed or written on various materials, often on durable blocks of stone or sheets of papyrus, which have been preserved primarily in drier climates.
@@
This traditional division within textual criticism—between the objective, technical analysis of manuscript data and the subjective, intuitive insight of an astute critic—is undergoing a significant transformation. Machine learning has the potential to contribute across all areas of philology, including those traditionally characterized by interpretation, conjecture, and creativity. Language models can be pretrained on the entire surviving textual record of ancient civilizations and fine-tuned to parse, classify, and analyze various aspects of words, phrases, and texts in these languages. They can be applied to explore relationships between manuscripts, identify prosodic, rhetorical, and narrative patterns, detect intertextual connections, recognize themes, analyze sentiments, summarize ideas, and translate texts (see review in [2]). More directly relevant to our purposes, they can be trained to date and geographically attribute texts, detect errors, generate corrections, and fill in missing words and phrases ([3, 4, 5, 6, 7]).
In the field of Greek documentary texts, Assael and colleagues [4] were and remain the leading pioneers. They trained a transformer model with a sparse attention mechanism (BigBird), combining word- and character-level tokenization, to date, attribute geographically, and reconstruct damaged sequences in ancient Greek inscriptions. On their test set of 7,811 inscriptions, they reported the following results:
@@
The aim of this experiment is to advance research in machine learning and textual criticism in three specific respects:
@@
Second, given the rapid advancements in deep learning and natural language processing over the last two years, it is worth exploring how newer, larger pretrained causal language models—such as Meta’s LLaMA, Google’s Gemma, Microsoft’s Phi, or OpenAI’s GPT-4o—can assist textual critics. One promising approach could involve enabling masked language modeling in these models and retraining them to fill in missing sequences of text within a bidirectional context (see the first step of the recipe developed by [8], LLM2Vec). However, in this preliminary study, I have opted for a simpler approach that is immediately applicable and open to experimentation with larger and newer models in the future: using a straightforward instruct template, prompting the model to generate the date, location, or reconstruction at the end of the sequence after a special token.
@@ 2 Method
Dates for both papyri and inscriptions were formatted as discrete numbers corresponding to the midpoint of their respective date spans. For dates with only a terminus post quem, they were set 25 years after the given date, while those with only a terminus ante quem were set 25 years earlier. Place name attributions for the papyri were normalized semi-manually due to inconsistencies across database entries. While more detailed processing would be beneficial in the future, the current approach – though admittedly rough (see https://github.com/ericu9500/PapyriAndInscriptions/train_data/data/places_and_dates.tsv) – was deemed sufficient for the proof-of-concept purposes of this study. All text editions were included in two versions: one in which philologists’ reconstructions of lost or uncertain characters were integrated, and another where these characters were replaced with “-”. Individual missing characters were represented by a single “-”, while longer stretches of missing text were indicated by 10 consecutive hyphens.
@@
In an additional step, the models were fine-tuned for text restoration on texts truncated to 50–750 characters to allow for comparison with the model developed by Assael and colleagues [4]. They were tested on a subset of my test set that coincided with theirs (entries with a PHI ID ending in 3). This subset included 264 samples with dates, 435 samples with places, and 4,110 masked samples based on 411 papyri for restoration (100 samples of each character length). Additionally, the models were evaluated on 2 previously unedited Uppsala papyri (1,620 masked samples) and 6 recently edited inscriptions (4,111 masked samples) to assess their performance on texts that could not have been included in the pretraining data of Meta’s base model. Different subpages of both http://papyri.info and https://inscriptions.packhum.org sporadically appear in the indices of the Common Crawl corpus (http://index.commoncrawl.org).
@@ 3.1 Calibration
  • Reconstruction: Three epochs for inscriptions (CER 27.9%, top-1 accuracy 52.2%, and top-20 accuracy 71.0%) and four epochs for papyri (CER 20.6%, top-1 accuracy 64.0%, and top-20 accuracy of 80.86%).
@@ 3.1.1 Continued training
For the dating task, performance plateaued after 1 additional epoch (for a total of 4 epochs) for the inscriptions, with an average deviation of 37.1 years and a median deviation of 3 years from the actual dating span. Training on the papyrus dataset for dating also plateaued after 1 additional epoch (for a total of 4 epochs), with an average deviation of 21.6 years and a median deviation of 0 years from the actual dating span.
@@ 3.1.2 Re-merging
3.1.3 (Imperfect) comparison with Ithaca
An accurate comparison with the state-of-the-art model for these tasks developed by Assael and colleagues [4] – Ithaca – is not possible, since the present model was trained for experimental purposes with a 95%/5% split as opposed to Ithaca’s 80/10/10. Yet some form of comparison is still useful. Accordingly, in a final step, the models were finetuned on the dataset where every entry was truncated to 50–750 characters (minimum and maximum length of entries for Ithaca). The resulting inscription restoration model and Ithaca were then run on two test sets: one solely derived from entries within my 5% test set that coincided with Assael and colleagues’ test split (where PHI ID ends in 3), and one derived from inscriptions that were not in the PHI or openly available on the internet as of December 2023 (4111 samples based on 6 inscriptions; see appendix). On the shared test set, the finetuned Llama model had an average CER of 22.55% (vs. Ithaca’s 27.0%), a top-1 accuracy of 61.4% (vs. 60.9%), and top-20 accuracy of 77.00% (vs. 71.6%). On the new inscriptions, it had a CER of 30.8% (vs. Ithaca’s 31.1%), a top-1 accuracy of 48.9% (vs. 46.5%), and a top-20 accuracy of 72.9% (vs. 68.3%).
@@
These further-tuned models were also tested on the entire 5% test set with good results. The inscription model had a final CER of 20.5%, a top-1 accuracy of 63.7%, and top-20 accuracy of 83.0%. The Papyrus model had a CER of 14.9%, top-1 accuracy of 73.5%, and top-20 accuracy of 85.9%. Accordingly they were chosen as the final models for restoration.
@@ 4 Discussion
Epigr_2_Llama3.1-8b-Instruct on my test 20.5 63.7 83.0 75.0 83.7 37.1
@@
Epigr_1_Llama3.1-8b-Instruct on shared test 22.6 61.4 81.6 65.0 78.4 55.7
Assael et al. (2022) on shared test set 26.3 61.8 78.3 69.0 80.2 48.3
Epigr_1_Llama3.1-8b-Instruct on new inscr. 30.8 48.9 72.9 - - -
@@
Papy_1_Llama3.1-8b-Instruct on my test 14.9 73.5 85.9 66.5 80.0 21.7
Papy_1_Llama3.1-8b-Instruct on new papyri 29.67 73.0 89.0 - - -
@@
Table 1 summarizes the results, providing a preliminary comparison between the models developed in this study and Ithaca [4]. While a new experiment fine-tuning LLaMA using an 80/10/10 split would be necessary for a more accurate comparison, the superior results on the test set with new inscriptions suggest that the models developed here generalize well to unseen data, rather than merely overfitting to the test set. Furthermore, the results on unedited or recently edited papyri and inscriptions imply that the strength of these models does not simply stem from possible exposure to test texts included in the data used during Meta’s pretraining process, but rather from the effects of the fine-tuning process itself. Although they exhibit a higher CER and lower top-1 and top-20 accuracy on this material, the fact that Ithaca’s performance was similarly affected, along with the variation in results between the two papyri, suggests that the challenge lies in the task itself rather than a lack of prior exposure.
In any case, the primary goal of this experiment was to assess whether a simple, instruction-based approach could compete with more established methods in this area, and the findings indicate that it can. These results justify further exploration through additional experiments involving smaller, larger, and more advanced base models, as well as alternative strategies for dataset formatting, masking, training, and model merging. The tools developed in this study perform at least comparably to the state-of-the-art in text restoration, demonstrating significant potential. However, the method’s effectiveness for dating and geographic attribution was less conclusive.
Moreover, these models offer a notable advantage: they are particularly useful in cases where scholars have an approximate sense of the number of missing characters but lack precise information about missing word boundaries—a frequent challenge in the study of ancient textual artifacts. Additionally, the papyri models introduced here represent a new resource for this material.
@@ 5 Conclusion
This experiment demonstrated the efficacy of fine-tuning pretrained causal language models for text restoration in ancient Greek inscriptions and documentary papyri using a simple instruction template. The method’s success in dating and geographic attribution was less definitive. While the results are preliminary, they suggest that this straightforward approach not only approaches but, in some respects, may surpass the current state of the art. The simplicity and scalability of this method, particularly the ease of fine-tuning newer models on the same data with minimal modification, is a key advantage. There is significant potential for further improvement through refined data cleaning, data augmentation, and masking strategies, optimized training parameters, and experimentation with smaller, larger, and more advanced base models as they become available.

### v3

# Instruction-Tuning Pretrained Causal Language Models for Text Restoration of Ancient Greek Papyri and Inscriptions

@@ Abstract
This article presents an experiment in fine-tuning a pretrained causal language model (Meta’s Llama 3.1 8B Instruct) to assist with restoring missing or illegible characters in ancient Greek inscriptions and documentary papyri.
Utilizing a straightforward instruction-based approach and a 95%/5% train/test split, the papyrus restoration model achieved a character error rate (CER) of 14.9%, a top-1 accuracy of 73.5%, and a top-20 accuracy of 86.0% for sequences up to 10 characters. A model was also fine-tuned for geographic attribution, reaching a top-1 accuracy of 66.4% and a top-3 accuracy of 79.9%. In chronological attribution, it demonstrated an average deviation of 21.7 years from the actual terminus post/ante quem, with a median deviation of 0 years.
For inscriptions, the restoration model achieved a CER of 20.5%, a top-1 accuracy of 63.7%, and a top-20 accuracy of 83.0% for sequences up to 10 characters. In geographic attribution, it attained a top-1 accuracy of 75.0% and a top-3 accuracy of 83.7%, while in dating, it had an average deviation of 37.1 years and a median deviation of 3 years from the actual date range.
Benchmarked against the state-of-the-art model (Ithaca) on a shared test set and on recently edited inscriptions, the instruction-tuned models excelled in text restoration, while also offering the practical advantage of ignoring spaces during reconstruction, which aligns with the scriptio continua of ancient textual artifacts. However, their performance in geographic and chronological attribution was lower than Ithaca’s. To evaluate the approach in a more even setup, the instruction model was retrained with an 80%/10%/10% train-validation-test split, and still outperformed Ithaca in text restoration.
The results are preliminary but suggest that fine-tuning larger pretrained causal language models using instruction templates for emendations and conjectures to ancient texts holds significant promise.
@@ 1 Introduction
A significant number of documentary texts in ancient Greek—such as announcements, laws, receipts, contracts, and letters—have survived from the Archaic period through to the early Byzantine era. These documents were inscribed or written on various materials, including blocks of stone and sheets of papyrus.
@@
This traditional division within textual criticism—between the objective, technical analysis of manuscript data and the subjective, intuitive insight of an astute critic—is undergoing a significant transformation. Machine learning has the potential to contribute across all areas of philology, including those traditionally characterized by interpretation, conjecture, and creativity. Language models can be pretrained on the entire surviving textual record of ancient civilizations and fine-tuned to parse, classify, and analyze various aspects of words, phrases, and texts in these languages. They can be applied to explore relationships between manuscripts, identify prosodic, rhetorical, and narrative patterns, detect intertextual connections, recognize themes, classify sentiments, summarize ideas, and translate texts (see review in [2]). More directly relevant to our purposes, they can be trained to date and geographically attribute texts, detect errors, generate corrections, and fill in missing words and phrases ([3, 4, 5, 6, 7]).
In the field of Greek documentary texts, Assael and colleagues [4] were the leading pioneers. They trained a transformer model with a sparse attention mechanism (BigBird), combining word- and character-level tokenization, to date, attribute geographically, and reconstruct damaged sequences in ancient Greek inscriptions. On their test set of 7,811 inscriptions, they reported the following results:
@@
The aim of the following experiment is to build on this work and advance research in machine learning and textual criticism in three specific respects:
@@
Second, given the rapid advancements in deep learning and natural language processing over the last two years, it is worth exploring how newer, larger pretrained causal language models—such as Meta’s LLaMA, Google’s Gemma, Microsoft’s Phi, or OpenAI’s GPT-4o—can assist textual critics. One promising approach could involve enabling masked language modeling in these models and retraining them to fill in missing sequences of text within a bidirectional context (see the first step of the recipe developed by [8], LLM2Vec). However, in this preliminary study, I have opted for a simpler approach that is immediately applicable and open to experimentation with larger and newer models in the future: using a straightforward instruction template, prompting the model to generate the date, location, or reconstruction at the end of the sequence after a special token.
@@ 2 Method
Dates for both papyri and inscriptions were formatted as discrete numbers corresponding to the midpoint of their respective date spans. For dates with only a terminus post quem, they were set 25 years after the given date, while those with only a terminus ante quem were set 25 years earlier. Place name attributions for the papyri were normalized semi-manually due to inconsistencies across database entries. While more detailed processing would be beneficial in the future, the current approach—though admittedly rough (see https://github.com/ericu9500/PapyriAndInscriptions/train_data/data/places_and_dates.tsv)—was deemed sufficient for the proof-of-concept purposes of this study. All text editions were included in two versions: one in which philologists’ reconstructions of lost or uncertain characters were integrated, and another where these characters were replaced with “-”. Individual missing characters were represented by a single “-”, while longer stretches of missing text were indicated by 10 consecutive hyphens.
@@
In an additional step, the models were fine-tuned for restoration on texts truncated to 50–750 characters to allow for comparison with the model developed by Assael and colleagues [4]. They were tested on a subset of my test set that coincided with theirs (entries with a PHI ID ending in 3). This subset included 264 samples with dates, 435 samples with places, and 4,110 masked samples based on 411 papyri for restoration (100 samples of each character length). Additionally, the models were evaluated on 2 previously unedited Uppsala papyri (1,620 masked samples) and 6 recently edited inscriptions (4,111 masked samples) to assess their performance on texts that could not have been included in the pretraining data of Meta’s base model. Subpages of both http://papyri.info and https://inscriptions.packhum.org sporadically appear in the indices of the Common Crawl corpus (http://index.commoncrawl.org). In a final step, the inscription restoration model was fine-tuned from scratch based on the papyrus restoration model using the same 80%/10%/10% train-validation-test split as Assael and colleagues [4], augmenting the training data with OpenAI’s GPT-4o-mini (see the dataset on https://huggingface.co/datasets/Ericu950/Inscriptions_2).
@@ 3.1 Calibration
  • Reconstruction: Three epochs for inscriptions (CER 27.9%, top-1 accuracy 52.2%, and top-20 accuracy 71.0%) and four epochs for papyri (CER 20.6%, top-1 accuracy 64.0%, and top-20 accuracy 80.86%).
@@ 3.1.1 Continued training
For the dating task, performance plateaued after 1 additional epoch (for a total of 4 epochs) for the inscriptions, with an average deviation of 37.1 years and a median deviation of 3 years from the actual dating span. Training on the papyrus dataset for dating also plateaued after 1 additional epoch (for a total of 4 epochs), with an average deviation of 21.7 years and a median deviation of 0 years from the actual dating span.
@@ 3.1.2 Re-merging
3.1.3 Full comparison with Ithaca
An accurate comparison with the state-of-the-art model for these tasks developed by Assael and colleagues [4] – Ithaca – was at first not possible, since the present model was trained for experimental purposes with a 95%/5% split as opposed to Ithaca’s 80%/10%/10%. Yet some form of comparison was considered useful. Accordingly, in a penultimate step, the models were fine-tuned on the dataset where every entry was truncated to 50–750 characters (minimum and maximum length of entries for Ithaca). The resulting inscription restoration model and Ithaca were then run on two test sets: one solely derived from entries within my 5% test set that coincided with Assael and colleagues’ test split (where PHI ID ends in 3), and one derived from inscriptions that were not in the PHI or openly available on the internet as of December 2023 (4111 samples based on 6 inscriptions; see appendix). On the shared test set, the fine-tuned Llama model had an average CER of 22.55% (vs. Ithaca’s 27.0%), a top-1 accuracy of 61.4% (vs. 60.9%), and top-20 accuracy of 77.00% (vs. 71.6%). On the new inscriptions, it had a CER of 30.8% (vs. Ithaca’s 31.1%), a top-1 accuracy of 48.9% (vs. 46.5%), and a top-20 accuracy of 72.9% (vs. 68.3%).
@@ 3.1.3 Full comparison with Ithaca
These further-tuned models were also tested on the entire 5% test set with good results. The inscription model had a final CER of 20.5%, a top-1 accuracy of 63.7%, and top-20 accuracy of 83.0%. The Papyrus model had a CER of 14.9%, top-1 accuracy of 73.5%, and top-20 accuracy of 85.9%.
In order to conduct a more exact comparison between the approach explored in this paper and Ithaca, an attempt to use the same 80%/10%/10% split was made. The papyrus restoration model was used as the base model. Only inscriptions with an ID not ending in 3 or 4 were used. For data augmentation, each inscription was processed with OpenAI’s GPT-4o mini to create a modified version of the text with 10 slight variations of the following system prompt:
Generate a modified version of this inscription for training a model to restore gaps in ancient inscriptions. Apply advanced data augmentation techniques, such as reordering words and sentences, and replacing as many words and phrases as possible with their synonyms, while strictly preserving Greek morphology, syntax, proper names, and place names. Ensure creativity in restructuring, but avoid altering the essential meaning. Only provide the altered inscription, without additional commentary.
Performance on the validation set plateaued after 8 epochs. On the exact same subsection of the data used in the previous test, it achieved a somewhat higher yet competitive CER of 21.8%, top-1 accuracy of 62.3%, and top-20 accuracy of 78.0%. On the test set of recently edited inscriptions, it outperformed the previous model with a CER of 28.6%, a top-1 accuracy of 51.3%, and top-20 accuracy of 68.6%.
@@ 4 Discussion
Epigr_1/2_Llama3.1-8b-Instruct on my test 20.5 63.7 83.0 75.0 83.0 37.1
@@
Epigr_3_Llama3.1-8b-Instruct on shared test 21.8 62.3 78.0 65.0 78.4 55.7
Assael et al. (2022) on shared test set 27.0 60.9 71.6 69.0 80.2 48.3
Epigr_3_Llama3.1-8b-Instruct on new inscr. 28.6 51.3 68.6 - - -
@@
Papy_1_Llama3.1-8b-Instruct on my test 14.9 73.5 85.9 66.4 79.9 21.7
Papy_1_Llama3.1-8b-Instruct on two new papyri 21.0/38.6 62.5/37.2 79.0/61.2 - - -
@@
Table 1 summarizes the results, providing a direct comparison between the models developed in this study and Ithaca [4]. The superior results on the shared test set as well as the test set with unedited or recently edited papyri and inscriptions suggest that the strength of these models does not simply stem from possible exposure to test texts included in the data used during Meta’s pretraining process, but rather from the effects of the fine-tuning process itself. Although they exhibit a higher CER and lower top-1 and top-20 accuracy on this material, the fact that Ithaca’s performance was similarly affected, along with the variation in results between the two different papyri, suggests that the challenge lies in the task itself rather than a lack of prior exposure.
In any case, the primary goal of this experiment was to assess whether a simple, instruction-based approach could compete with more established methods in this area, and the findings indicate that it can. These results justify further exploration through additional experiments involving smaller, larger, and more advanced base models, as well as alternative strategies for dataset formatting, masking, training, and model merging. The tools developed in this study outperform the state-of-the-art in text restoration, demonstrating significant potential. However, the method’s effectiveness for dating and geographic attribution was less conclusive.
Moreover, these models offer a notable advantage: they are particularly useful in cases where scholars have an approximate sense of the number of missing characters but lack precise information about missing word boundaries—a frequent challenge in the study of ancient textual artifacts. Furthermore, the study proved the effectiveness of using large language models for data augmentation to improve text restoration models for Ancient Greek texts. Additionally, the papyri models introduced here represent a completely new resource for this material.
@@ 5 Conclusion
This experiment demonstrated the efficacy of fine-tuning pretrained causal language models for text restoration in ancient Greek inscriptions and documentary papyri using a simple instruction template, surpassing the current state of the art. The simplicity and scalability of this method, particularly the ease of fine-tuning newer models on the same data with minimal modification, is a key advantage. There is significant potential for further improvement through refined data cleaning, data augmentation, and masking strategies, optimized training parameters, and experimentation with smaller, larger, and more advanced base models as they become available. However, the method was less successful than previous models for dating and geographic attribution.